OpenStudy (skarlettrose186):

Which list of ordered pairs represents solutions to x + y = 2 ? (-4, 6), (0, 2), (4, 2) (-4, -6), (0, 2), (4, 2) (-4, 6), (0, 2), (4, -2)

jhonyy9 (jhonyy9):

x+y=2 what is the general form of this equation y=f(x) = ?

jhonyy9 (jhonyy9):

given x+y=2 so from this y= ?

OpenStudy (skarlettrose186):

y=2-x

jhonyy9 (jhonyy9):

ok. i will tel you y= -x+2 yes ? so now just give values to x and calculi the y values what will result

jhonyy9 (jhonyy9):

or you check these given parentheses for what will get right values - ok. ?

OpenStudy (skarlettrose186):

(-4, -6), (0, 2), (4, 2) is this the right answer

jhonyy9 (jhonyy9):

in case of this so mean that for x = -4 you need getting y=-6 do you can check it this ?

jhonyy9 (jhonyy9):

why you wrote -6 when there is 6 with -4 ?

OpenStudy (skarlettrose186):

(-4, 6), (0, 2), (4, -2) is this right

jhonyy9 (jhonyy9):

than you try y= -x+2 for x= -4 you will get y= 6 ?

OpenStudy (skarlettrose186):

no

jhonyy9 (jhonyy9):

why no ? just substitute in place of x the -4 and make the calculi so what will get ?

OpenStudy (skarlettrose186):

i'm confused

OpenStudy (skarlettrose186):

is it b or c

jhonyy9 (jhonyy9):

why ? this is very very easy y=-x+2 6 = -(-4) +2 this is true or false ?

OpenStudy (skarlettrose186):

true

jhonyy9 (jhonyy9):

yes this is true ?

OpenStudy (skarlettrose186):

so its (-4, 6), (0, 2), (4, -2) right

jhonyy9 (jhonyy9):

yes for (-4,6) is true than now you need check again for x=0 you need getting for y=2 yes ? and for x=4 need getting y=-2 so than these will get right too this mean that the right answer is choice a. ok. ?

OpenStudy (skarlettrose186):

ok thanks

jhonyy9 (jhonyy9):

yw. was my pleasure good luck bye bye